# Connections between Combinations Without Specified Separations and Strongly Restricted Permutations, Compositions, and Bit Strings

### v2

@@ 3 Bijections between strongly restricted permutations and restricted combinations
Figure 5: A 13-board tiled with an up $(\frac{1}{2},9)$-fence ($F$), three down $(\frac{1}{2},2)$-fences ($\bar{F}$), and 9 squares.
@@
Let $P^{\mathcal{D}}_{n}$ denote the number of permutations $\pi$ of $\mathbb{N}_{n}$ such that $\pi(i)-i\in\mathcal{D}$ for all $i\in\mathbb{N}$ and $P^{\mathcal{D}}_{n,k}$ denote the number of such permutations with $k$ excedances. Then for $m,j\geq 1$, $i,k\geq 0$, and $r=0,\ldots,m-1$, we have
@@
where $f^{(t)}_{n}$ is the $n$-th $(1,t)$-bonacci number defined by $f^{(t)}_{n}=f^{(t)}_{n-1}+f^{(t)}_{n-t}+\delta_{n,0}$, $f^{(t)}_{n<0}=0$, the $(1,t)$-bonacci polynomial $f^{(t)}_{n}(x)$ is defined by $f^{(t)}_{n}(x)=f^{(t)}_{n-1}(x)+xf^{(t)}_{n-t}(x)+\delta_{n,0}$, $f^{(t)}_{n<0}(x)=0$, and $[x^{k}]R(x)$ denotes the coefficient of $x^{k}$ in $R(x)$.
@@ 4 Connections with compositions
$\mathcal{Q}=\{0,1,\ldots,l-1,l+g,l+g+1,\ldots,2l+g-1,2(l+g),2(l+g)+1,\ldots,\\ (t-1)(l+g),\ldots,(t-1)(l+g)+r-1=q>0\}$
for some $l\geq g>0$, $t\in\mathbb{Z}^{+}$, and, for $t>1$, $r\geq(t-1)(l+g)-1$, then $S_{n}^{\mathcal{Q}}$ is the number of compositions of $n+q$ into parts 1 and $q+1$ when $t=1$ and parts 1, $q+1$, $q+l+1,\ldots,q+l+g$, $q+2l+g+1,\ldots,q+2(l+g),\ldots,q+(t-1)l+(t-2)g+1,\ldots,q+(t-1)(l+g)$ otherwise. Note that $\mathcal{Q}$ reduces to $\{0,1,\ldots,r-1=q\}$ when $t=1$.
@@
Four classes of $\mathcal{Q}$ whose associated digraphs possess common nodes have been investigated (Theorems 5–8 in [2]). One of these classes (Theorem 8 in [2]) always has at least two inner cycles; the simplest member of this class is $\mathcal{Q}=\{1,4\}$ (see Fig. 3). Another class (Theorem 7 in [2]) has just one inner cycle but always has the metatiles $C_{[q+1]}C_{[l]}^{2}$ and $C_{[q+1]}S^{a-1}C_{[2l]}S$, where $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$, which are of the same length (see Fig. 5 in [2]). The metatile-generating digraph of one of the remaining classes has $a$ inner cycles (Theorem 5 in [2]). The $a=1$ case leads to the following result.
@@
The following theorem originates from Theorem 6 in [2]. The OEIS sequences to which it applies along with the corresponding $\mathcal{Q}$ are as follows: A224809 {2,4}, A375981 {1,4,5}, A375982 {2,3,5}, A375983 {2,4,5}.
@@
Let $\theta$ be the bit string representation of $\mathcal{Q}$ whereby the $j$-th bit from the right of $\theta$ is 1 if and only if $j\in\mathcal{Q}$. By $\lfloor\theta/2^{b}\rfloor$ we mean discarding the rightmost $b$ bits in $\theta$ and shifting the remaining bits to the right $b$ places. Let $p_{i}\in\mathbb{N}_{q}-\mathcal{Q}$ for $i=1,\ldots,a$ and $p_{i}<p_{i+1}$. Using $\mid$ to denote the bitwise OR operation, if $\theta\mid\lfloor\theta/2^{p_{i}-1}\rfloor$ for each $i=1,\ldots,a-1$ is all ones after discarding the leading zeros, $a\geq 2$, and $p_{a}=q-r$ (which implies that $r\geq 1$), then if (a) $q=2r+1$ or (b) $q>2r+1$ and $1\leq p_{a-1}\leq r$, then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts 1, $q+1$, $q+1+p_{i}$ for $i=1,\ldots,a-1$, and $q+1+j(q-r)$ for $j\in\mathbb{Z}^{+}$.
@@ 5 Connections with bit strings
Others have noted that the number of length-$n$ bit strings (or binary words) having no two 1s that have positions in the string that differ by $q$ (or, equivalently, are separated from one another by $q-1$ digits) is $S^{\{q\}}_{n}$ [14]. This can be generalized by considering the following bit string $s$ representing subset $\mathcal{S}$: the $j$-th bit from the right of $s$ is 1 if and only if $j\in\mathcal{S}$. (This representation is used in the program in the Appendix.) Then it is clear that $S^{\mathcal{Q}}_{n,k}$ is the number of length-$n$ bit strings that contain $k$ 1s placed in such a way that the difference in positions of any two 1s does not equal an element of $\mathcal{Q}$. We now mention an instance of this that corresponds to A130137 in the OEIS [20]. If one instead places the same restrictions on 0s rather than 1s, for the $\mathcal{Q}=\{1\}$ case, the bit strings (i.e., those containing no 00) are called Fibonacci binary words, so named as the number of such words of length $n$ is $F_{n+2}$. A130137 is described as the number of length-$n$ Fibonacci binary words that do not contain the substring 0110. This is $S^{\{1,3\}}_{n}$ since the three other possible disallowed substrings that only prevent the occurrence of two 0s whose positions differ by 3 (i.e., $0$, 0010, and 0100) do not need to be mentioned; they each contain 00 and so cannot be present in Fibonacci binary words.
@@
In the rest of this section we show that the elements of $\mathcal{Q}$ concerned with the bijection given in Theorem 20 are always a well-based sequence and hence the generating function for the number of equivalence classes and the recursion relation for the number of equivalence classes whose set representation is of size $k$ are straightforward to obtain. The sequence $q_{1},q_{2},\ldots,q_{m}$, where $q_{i}<q_{i+1}$ for all $i$, is said to be a well-based sequence if $q_{1}=1$ and for all $j>1$ and $\Delta=1,\ldots,q_{j}-1$, there is some $q_{i}$ such that $q_{j}=q_{i}+\Delta$ [12, 21]. In the present context, we can regard $q_{i}$ as being the elements of $\mathcal{Q}$ with $q_{m}=q$, its largest element. In the proof of Theorem 21, we will use the following definition which is equivalent aside from also classing the empty set as being well based: the sequence is well based if $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$ is zero or if, for all $i,j=1,\ldots,a$ (where $i$ and $j$ can be equal), $p_{i}+p_{j}\notin\mathcal{Q}$, where the $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$. E.g., the only well-based sequences of length 3 are the elements of the sets $\{1,2,3\}$, $\{1,2,4\}$, $\{1,2,5\}$, and $\{1,3,5\}$.
@@
The generating function $\tilde{g}(x,y)$, whose coefficient of $x^{n}y^{k}$ equals $E^{(\omega)}_{n,k}$, the number of equivalence classes of length-$n$ binary words with respect to a length-$l$ subword $\omega$ that have a set representation of size $k$, is given by
@@
where $\bar{c}=\sum_{i=1}^{a}x^{p_{i}}y$, $a=\left\lvert\mathbb{N}_{q}-\mathcal{Q}\right\rvert$, $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}y$, the set $\mathcal{Q}$ is obtained from $\omega$ using the procedure in Theorem 20, $q_{i}$ are the elements of $\mathcal{Q}$ of which the largest is $q$ (with $q=0$ when $\mathcal{Q}=\{\}$), and $p_{i}$ are the elements of $\mathbb{N}_{q}-\mathcal{Q}$, provided that $q=l-1$.
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and the sums in (8) are omitted leaving $\tilde{g}(x,y)=1/(1-x-xy)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n,k}=\tbinom{n}{k}$ since this is the number of length-$n$ binary words that contain $k$ 1s. The generating function for $\tbinom{n}{k}$ is $1/(1-x-xy)$. For $q>0$, since $E^{(\omega)}_{n}=E^{(\omega)}_{n,0}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 20, $E^{(\omega)}_{n,k}=S^{\mathcal{Q}}_{n-q,k}$ for $n\geq q$ and from Theorem 1 $S^{\mathcal{Q}}_{n-q,k}=B^{\mathcal{Q}}_{n,k}$ for $n\geq q$, overall we have $E^{(\omega)}_{n,k}=B^{\mathcal{Q}}_{n,k}$. Hence $\tilde{g}(x,y)=G(x,y)$, the generating function such that the coefficient of $x^{n}y^{k}$ is the number of tilings of an $n$ board that contain $k$ combs. The digraph for tiling a board with squares and $\mathcal{Q}$-combs when the elements of $\mathcal{Q}$ are a well-based sequence is shown in Fig. 9 [2]. The digraph has a common node ($\sigma_{1}$) and the inner cycles ($S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a$) have lengths $p_{i}$. There is a single common circuit ($C_{[q+1]}S^{a}$), which is of length $q+1$, and one outer cycle ($S_{[1]}$), which is of length 1. The common circuit and all the inner cycles contain a single comb. From Theorem 5 we then have
@@ 6 Discussion
Figure 10: Digraph for restricted-overlap tiling with squares ($S$) and $(2,4,1)$-combs ($C$), which corresponds to the $\mathcal{Q}=\{1,6\}$ case.
Theorem 6, which we derived here in order to obtain the generating function for $S^{\{1,5\}}_{n}$, will be used elsewhere to obtain generating functions for $S^{\mathcal{Q}}_{n,k}$ for various other classes of $\mathcal{Q}$. The metatile-generating digraph corresponding to the next case in the $\{1,q\}$ series (Fig. 10) does not fall into any of the classes of digraph that we have derived a generating function for so far. However, the systematic application of the digraph generating function rules (Lemmas 2, 3, and 4) can still be used to yield the generating function in this instance.
Appendix: C program for finding $S_{n}$ and $S_{n,k}$
@@
// anything after // on the same line is a comment
// program name: rcl.c (restricted combinations on a line)
@@
// to create an executable called rcl using the GNU C compiler enter
// gcc rcl.c -o rcl
@@
// the variable type unsigned int typically uses 4 bytes
// hence MAXn is typically 32
#define MAXn 8*sizeof(unsigned int)
@@
unsigned int Q=0,s,ss; // bit strings giving Q, subset, and shifted subset
@@
printf("%d,",S[lastn][j]); // lastn now equals MAXn

### v3

@@ 3.2 Connection with the class $\mathcal{Q}=\{m,2m,\ldots,jm\}$
Figure 6: A 13-board tiled with an up $(\frac{1}{2},9)$-fence ($F$; blue), three down $(\frac{1}{2},2)$-fences ($\bar{F}$; yellow, orange, and ochre), and 9 squares.
@@
We let $P^{\mathcal{D}}_{n}$ denote the number of permutations $\pi$ of $\mathbb{N}_{n}$ such that $\pi(i)-i\in\mathcal{D}$ for all $i\in\mathbb{N}_{n}$ and $P^{\mathcal{D}}_{n,k}$ denote the number of such permutations with $k$ excedances. We now use the bijection and previously obtained results to relate $P^{\{-m,0,jm\}}_{n}$ and $P^{\{-m,0,jm\}}_{n,k}$ to some generalized Fibonacci numbers and the coefficients of their corresponding polynomials, respectively. We let $f^{(t)}_{n}$ denote the $n$-th $(1,t)$-bonacci number defined by $f^{(t)}_{n}=f^{(t)}_{n-1}+f^{(t)}_{n-t}+\delta_{n,0}$, $f^{(t)}_{n<0}=0$, and $f^{(t)}_{n}(x)$ denote the $(1,t)$-bonacci polynomial defined by $f^{(t)}_{n}(x)=f^{(t)}_{n-1}(x)+xf^{(t)}_{n-t}(x)+\delta_{n,0}$, $f^{(t)}_{n<0}(x)=0$.
@@
For $m,j\geq 1$, $i,k\geq 0$, and $r=0,\ldots,m-1$, we have
@@
where $[x^{k}]R(x)$ denotes the coefficient of $x^{k}$ in $R(x)$.
@@ 4 Connections with compositions
$\mathcal{Q}=\{0,1,\ldots,l-1,l+g,l+g+1,\ldots,2l+g-1,2(l+g),2(l+g)+1,\ldots,\\ (t-1)(l+g),\ldots,(t-1)(l+g)+r-1\equiv q\}$
for some $l\geq g>0$, $t\in\mathbb{Z}^{+}$, $q>0$, and, for $t>1$, $r\geq(t-1)(l+g)-1$, then $S_{n}^{\mathcal{Q}}$ is the number of compositions of $n+q$ into parts 1 and $q+1$ when $t=1$ and parts 1, $q+1$, $q+l+1,\ldots,q+l+g$, $q+2l+g+1,\ldots,q+2(l+g),\ldots,q+(t-1)l+(t-2)g+1,\ldots,q+(t-1)(l+g)$ otherwise. Note that $\mathcal{Q}$ reduces to $\{0,1,\ldots,r-1=q\}$ when $t=1$.
@@
Four classes of $\mathcal{Q}$ whose associated digraphs possess common nodes have been investigated (Theorems 5–8 in [2]). One of these classes (Theorem 8 in [2]) always has at least two inner cycles; the simplest member of this class is $\mathcal{Q}=\{1,4\}$ (see Fig. 3). Another class (Theorem 7 in [2]) has just one inner cycle but always has the metatiles $C_{[q+1]}C_{[l]}^{2}$ and $C_{[q+1]}S^{a-1}C_{[2l]}S$, where $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$, which are of the same length (see Fig. 5 in [2]). The metatile-generating digraph of one of the remaining classes has $a$ inner cycles (Theorem 5 in [2]). The $a=1$ case leads to the following result.
@@
The following theorem originates from Theorem 6 in [2] and as it is a rather general theorem, it employs some extra notation that we now describe. We let $\theta$ be the bit string representation of $\mathcal{Q}$ whereby the $j$-th bit from the right in $\theta$ is 1 if and only if $j\in\mathcal{Q}$. By $\lfloor\theta/2^{b}\rfloor$ we mean discarding the rightmost $b$ bits in $\theta$ and shifting the remaining bits to the right $b$ places. We use $\mid$ to denote the bitwise OR operation.
@@
Let $p_{i}\in\mathbb{N}_{q}\setminus\mathcal{Q}$ for $i=1,\ldots,a$ and $p_{i}<p_{i+1}$. If $\theta\mid\lfloor\theta/2^{p_{i}-1}\rfloor$ for each $i=1,\ldots,a-1$ is all ones after discarding the leading zeros, $a\geq 2$, and $p_{a}=q-r$ (which implies that $r\geq 1$), then if (a) $q=2r+1$ or (b) $q>2r+1$ and $1\leq p_{a-1}\leq r$, then $S^{\mathcal{Q}}_{n}$ is the number of compositions of $n+q$ into parts 1, $q+1$, $q+1+p_{i}$ for $i=1,\ldots,a-1$, and $q+1+j(q-r)$ for $j\in\mathbb{Z}^{+}$.
@@
The OEIS sequences to which the above theorem applies along with the corresponding $\mathcal{Q}$ are as follows: A224809 {2,4}, A375981 {1,4,5}, A375982 {2,3,5}, A375983 {2,4,5}.
@@ 5 Connections with bit strings
Others have noted that the number of length-$n$ bit strings (or binary words) having no two 1s that have positions in the string that differ by $q$ (or, equivalently, are separated from one another by $q-1$ digits) is $S^{\{q\}}_{n}$ [14]. This can be generalized by considering the following bit string $s$ representing subset $\mathcal{S}$: the $j$-th bit from the right in $s$ is 1 if and only if $j\in\mathcal{S}$. (This representation is used in the program in the Appendix.) Then it is clear that $S^{\mathcal{Q}}_{n,k}$ is the number of length-$n$ bit strings that contain $k$ 1s placed in such a way that the difference in positions of any two 1s does not equal an element of $\mathcal{Q}$. We now mention an instance of this that corresponds to A130137 in the OEIS [20]. If one instead places the same restrictions on 0s rather than 1s, for the $\mathcal{Q}=\{1\}$ case, the bit strings (i.e., those containing no 00) are called Fibonacci binary words, so named as the number of such words of length $n$ is $F_{n+2}$. A130137 is described as the number of length-$n$ Fibonacci binary words that do not contain the substring 0110. This is $S^{\{1,3\}}_{n}$ since the three other possible disallowed substrings that only prevent the occurrence of two 0s whose positions differ by 3 (i.e., $0$, 0010, and 0100) do not need to be mentioned; they each contain 00 and so cannot be present in Fibonacci binary words.
@@
In the rest of this section we show that the elements of $\mathcal{Q}$ concerned with the bijection given in Theorem 21 are always a well-based sequence and hence the generating function for the number of equivalence classes and the recursion relation for the number of equivalence classes whose set representation is of size $k$ are straightforward to obtain. The sequence $q_{1},q_{2},\ldots,q_{m}$, where $q_{i}<q_{i+1}$ for all $i$, is said to be a well-based sequence if $q_{1}=1$ and for all 2-partitions of $q_{j}$, for $j=2,\ldots,q_{m}$, at least one of the two parts is a $q_{i}$ [11, 22]. Thus, for example, $1,3,4$ is not a well-based sequence since $4=2+2$ and 2 is not in the sequence. In the present context, we can regard $q_{i}$ as being the elements of $\mathcal{Q}$ with $q_{m}=q$, its largest element. In the proof of Theorem 22, we will use the following definition which is equivalent aside from also classing the empty set as being well based: the sequence is well based if $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ is zero or if, for all $i,j=1,\ldots,a$ (where $i$ and $j$ can be equal), $p_{i}+p_{j}\notin\mathcal{Q}$, where the $p_{i}$ are the elements of $\mathbb{N}_{q}\setminus\mathcal{Q}$. E.g., the only well-based sequences of length 3 are the elements of the sets $\{1,2,3\}$, $\{1,2,4\}$, $\{1,2,5\}$, and $\{1,3,5\}$.
@@
The following corollary gives a generating function for $E^{(\omega)}_{n,k}$, the number of equivalence classes of length-$n$ binary words with respect to a length-$l$ subword $\omega$ that have a set representation of size $k$.
@@
The generating function $\tilde{g}(x,y)$, whose coefficient of $x^{n}y^{k}$ equals $E^{(\omega)}_{n,k}$, is given by
@@
where $\bar{c}=\sum_{i=1}^{a}x^{p_{i}}y$, $a=\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$, $c=1+\sum_{i=1}^{\left\lvert\mathcal{Q}\right\rvert}x^{q_{i}}y$, the set $\mathcal{Q}$ is obtained from $\omega$ using the procedure in Theorem 21, $q_{i}$ are the elements of $\mathcal{Q}$ of which the largest is $q$ (with $q=0$ when $\mathcal{Q}=\{\}$), and $p_{i}$ are the elements of $\mathbb{N}_{q}\setminus\mathcal{Q}$, provided that $q=l-1$.
@@
We first deal with the $q=0$ case. Then $\mathcal{Q}=\{\}$ and so $\bar{c}=0$ and $c=1$ which leaves $\tilde{g}(x,y)=1/(1-x-xy)$. In this case, each binary word is in an equivalence class of its own and so $E^{(\omega)}_{n,k}=\tbinom{n}{k}$ since this is the number of length-$n$ binary words that contain $k$ 1s. The generating function for $\tbinom{n}{k}$ is $1/(1-x-xy)$. For $q>0$, since $E^{(\omega)}_{n}=E^{(\omega)}_{n,0}=1$ for $10\leq n\leq q-$ 1, and, by Theorem 21, $E^{(\omega)}_{n,k}=S^{\mathcal{Q}}_{n-q,k}$ for $n\geq q$ and from Theorem 1 $S^{\mathcal{Q}}_{n-q,k}=B^{\mathcal{Q}}_{n,k}$ for $n\geq q$, overall we have $E^{(\omega)}_{n,k}=B^{\mathcal{Q}}_{n,k}$. Hence $\tilde{g}(x,y)=G(x,y)$, the generating function such that the coefficient of $x^{n}y^{k}$ is the number of tilings of an $n$ board that contain $k$ combs. The digraph for tiling a board with squares and $\mathcal{Q}$-combs when the elements of $\mathcal{Q}$ are a well-based sequence is shown in Fig. 10 [2]. The digraph has a common node ($\sigma_{1}$) and the inner cycles ($S^{i-1}C_{[p_{i}]}$ for $i=1,\ldots,a$) have lengths $p_{i}$. There is a single common circuit ($C_{[q+1]}S^{a}$), which is of length $q+1$, and one outer cycle ($S_{[1]}$), which is of length 1. The common circuit and all the inner cycles contain a single comb. From Theorem 5 we then have
@@ 6 Discussion
Theorem 6, which we derived here in order to obtain the generating function for $S^{\{1,5\}}_{n}$, will be used elsewhere to obtain generating functions for $S^{\mathcal{Q}}_{n,k}$ for various other classes of $\mathcal{Q}$. As $\left\lvert\mathbb{N}_{q}\setminus\mathcal{Q}\right\rvert$ increases, the digraphs become too complicated for the application of general results analogous to Theorem 6. In such instances, the repeated use of Lemmas 2, 3, and 4 will yield the generating function, as was demonstrated in §3.1 for the $\mathcal{Q}=\{1,6\}$ case, although this becomes laborious for larger digraphs. Another approach is to employ a transfer matrix technique. For a digraph with $v$ nodes, the basic method gives the following expression for the generating function [21]:
$G(x,y)=\frac{\det(\mathbf{I}_{v-1}-\mathbf{\check{T}})}{\det(\mathbf{I}_{v}-\mathbf{T})},$ (13)
where $\mathbf{I}_{m}$ is the $m\times m$ identity matrix and the $(i,j)$-th entry of transfer matrix $\mathbf{T}$ is the generating function for the direct connection from the $j$-th node to the $i$-th node. For convenience, one insists that the 0 node is the first node. Then $\mathbf{\check{T}}$ is obtained from $\mathbf{T}$ by removing the first column and first row. For example, choosing the $10^{2}$ 1 and $01$ nodes to be, respectively, the second and third nodes, the transfer matrix for the digraph in Figure 3 is
$\mathbf{T}=\begin{pmatrix}x&0&1\\ x^{5}y&0&x^{3}y\\ 0&1+x^{2}y&0\end{pmatrix},$
and (13) gives $G(x,y)=(1-x^{3}y-x^{5}y^{2})/(1-x-x^{3}y+x^{4}y-x^{5}(y+y^{2})+x^{6}y^{2}-x^{7}y^{2})$. As expected, this reduces to the first term in parentheses in (4) on replacing $y$ by 1. More sophisticated techniques involve first reducing the dimensions of the transfer matrices used in an analogous formula to (13) [12].
Acknowledgements
The author thanks the editors and two anonymous referees for their time and the detailed and helpful comments. He also thanks Alex Sáiz for testing the program on his 64-bit machine.
Appendix: C program for finding $S^{\mathcal{Q}}_{n}$ and $S^{\mathcal{Q}}_{n,k}$
@@ Acknowledgements
// Anything after // on the same line is a comment.
// Program name: rcl.c (restricted combinations on a line).
@@
// To create an executable called rcl using the GNU C compiler enter:
// gcc rcl.c -O2 -o rcl
@@
// the variable type unsigned long int typically uses 4 bytes
// on 32-bit machines and 8 bytes on 64-bit machines
// hence MAXn is typically 32 or 64, respectively
#define MAXn 8*sizeof(unsigned long int)
@@
unsigned long int Q=0,s,ss; // bit strings giving Q, subset, and shifted subset
@@
printf("%ld,",S[lastn][j]); // lastn now equals MAXn